Find the standard deviation of the expected value of the probability distribution, X={1, 2, 3, 4, 5} P(X)={.10, .35, .15, .23, .17}


1
Expert's answer
2022-03-07T09:27:02-0500

To find the standard deviation of the probability distribution, we can use the following formula:


"\sigma=\sqrt{\sum(x_i-\mu)^2\cdot P(x_i)}"

where:

"x_i:" the "i^{th}" value

"\mu:" the mean of the distribution

"P(x_i):" the probability of the "i^{th}" value


"\mu=1\cdot0.10+2\cdot0.35+3\cdot0.15+4\cdot0.23+5\cdot0.17=3.02" "X-\mu=\begin{Bmatrix}
 1-3.02, 2-3.02, 3-3.02, 4-3.02, 5-3.02
\end{Bmatrix}="

"=\begin{Bmatrix}
-2.02, -1.02, -0.02, 0.98, 1.98
\end{Bmatrix}"

"\sigma=\sqrt{(-2.02)^2\cdot 0.10+(-1.02)^2\cdot 0.35+(-0.02)^2\cdot 0.15+0.98^2\cdot 0.23+1.98^2\cdot 0.17}="

"=\sqrt{1.6596}=1.288"